OpenStudy (anonymous):

What percent of 65 is 162.5?

OpenStudy (anonymous):

x% (65) = 162.5 x%= 162.5/65 x% = 2.5%

OpenStudy (anonymous):

here i forgot to x/100 = 2.5 x=250 so the answer is 250%